# PSZ2 G181.06+48.47 I: X-ray exploration of a low-mass cluster with exceptionally-distant radio relics

### v1

@@ Abstract
Relics are diffuse, highly-polarized radio sources that trace merger-driven shocks at the periphery of merging galaxy clusters. The LOFAR survey recently discovered a rare example of double relics in the low-mass cluster PSZ2 G181.06+48.47. Through a detailed exploration of new Chandra and XMM-Newton observations, we reveal that PSZ2 G181.06+48.47 has a lower mass ($M_{\rm 500,X}=2.57^{+0.37}_{-0.38}\times 10^{14}\;{\rm M_{\odot}}$) than previously thought. Despite its cool global temperature of $kT_{500}=3.86^{+0.27}_{-0.29}$ keV, PSZ2 G181.06+48.47 is one of the most disturbed clusters in the Planck sample, with a complex morphological and thermodynamic structure. We discover a set of three discontinuities within $<500$ kpc of the cluster center, and, from a surface brightness analysis, place $5\sigma$ upper limits of $\mathcal{M}_{NE}<1.43$ and $\mathcal{M}_{SW}<1.57$ for any shock associated with the relic locations. We also revise established scaling relations for double radio-relics by adding 12 new systems not included in previous work. The PSZ2 G181.06+48.47 relics have the widest separation (scaled for $r_{500}$) of all known double-relic systems. The exceptional distance from the cluster center ($>r_{200}$), indicates the relics may be associated with shocks in the “run-away” phase. We propose that this late-stage, post-apocenter merger is captured as the two subclusters with a mass ratio of 1.2–1.4 fall back into each other. The outer relic shocks were likely produced at the first core passage, while the inner discontinuities are associated with the second infall.
@@
facilities: XMM, CXO, LOFAR, SDSS, Pan-STARRSsoftware: pyproffit (Eckert et al., 2020), Astropy (Astropy Collaboration et al., 2013, 2018), APLpy (Robitaille & Bressert, 2012), Matplotlib (Hunter, 2007), TOPCAT (Taylor, 2005), DS9 (Joye & Mandel, 2003), CARTA (Comrie et al., 2021), CIAO (Fruscione et al., 2006), SAS
@@ 1 Introduction
Thermal and non-thermal components of the ICM such as shocks, cold fronts, turbulence, magnetic fields, and cosmic ray particles make clusters unique laboratories to study plasma physics (e.g. Sarazin, 1986, 2004; van Weeren et al., 2019). Thanks to the combination of sub-arcsecond angular resolution (0.5″, with Chandra) and sensitivity to diffuse extended thermal emission achievable by X-ray telescopes, the impact that merger events have on the plasma atmospheres and substructures can be directly mapped, for example, through sharp X-ray surface brightness edges, such as shocks and cold fronts (e.g. Markevitch et al., 2002, 2005; Markevitch & Vikhlinin, 2007; Owers et al., 2009; Ogrean & Brüggen, 2013; Eckert et al., 2016; Akamatsu et al., 2017; Zhang et al., 2021b; Diwanji et al., 2024). Unlike shocks, where the gas density and temperature are higher in the downstream compared to the upstream, cold fronts are formed at the boundary of gas clouds moving through a hotter and rarefied surrounding medium (see the review by Markevitch & Vikhlinin, 2007). Unlike shock fronts that are more strictly associated with cluster mergers, cold fronts have been detected in both merging and relaxed clusters (e.g., Botteon et al., 2018; Ghizzardi et al., 2010; Markevitch & Vikhlinin, 2007; Paterno-Mahler et al., 2013).
Figure 1: Multiwavelength view of PSZ2 G181.06+48.47, showing the central $15\arcmin\times 14\arcmin$ area. Left: 0.5–2 keV XMM-Newton image of the cluster, with LOFAR 140 MHz radio contours drawn $[1,2,4,8...]\times 3.0\sigma_{\rm rms}$, unveiling the cluster as a clear merger with an elongated, double-peaked X-ray morphology. Center: RGB Pan-STARRS image, with radio (green) and XMM-Newton (magenta, drawn at $[2,4,6,8...]\times 10^{-8}\,{\rm cts/s}$) contours overplotted. Saffron yellow squares mark spectroscopically-confirmed cluster members, including two BCGs roughly following the XMM-Newton X-ray peaks. The white plus marks the center of the system. Right: The 140 MHz LOFAR image at 6″ resolution reveals two spectacular diffuse radio sources on either side of the XMM-Newton X-ray detection of the cluster ICM (contours in magenta). The white, dashed circle represents $r_{\rm 500,SZ}=1.06$ Mpc.
@@
In rare cases, merging systems can exhibit two relics on opposite sides of the cluster. To our knowledge, only 30 examples of merging clusters hosting double radio relics have been discovered so far (de Gasperin et al., 2014; van Weeren et al., 2019; Botteon et al., 2020a; Jones et al., 2023). For example, only about 1/3 of simulated radio relic clusters may host candidate double relics of relatively similar luminosity (luminosity ratio $>0.3$, Lee et al., in prep). More importantly, double relics are more likely to be observed in simple binary cluster mergers with a favorable orientation close to the plane of the sky (Golovich et al., 2019), providing a unique laboratory to study the shocks with less interference by projection effects.
@@
While many X-ray shocks associated with radio relics have been successfully studied, shocks created in low-mass clusters with lower ICM sound speeds and, hence, slower shock velocities remain under-explored (Eckert et al., 2016; Akamatsu et al., 2017; Urdampilleta et al., 2018). Such shocks probe a key region of parameter space and are critical to constraining shock acceleration models. For example, standard diffusive shock acceleration (DSA, e.g. Bell, 1978; Drury, 1983; Blandford & Eichler, 1987), where particles from the thermal pool are directly accelerated by the merger-induced shock front, has been proposed as an effective mechanism for the origin of radio relics (for details see Brunetti & Jones, 2014). Direct (e.g. Hitomi Collaboration et al., 2018) and indirect (e.g. Dupourqué et al., 2024; Lovisari et al., 2024; Romero et al., 2024) X-ray measurements suggest average Mach numbers are below 2 in clusters. However, for such low Mach number shocks ($\mathcal{M}<2$; Ryu et al., 2019), DSA is not sufficient to explain the observed bright emission associated with large radio relics due to the low acceleration efficiency (e.g. Botteon et al., 2020a). Only a small fraction ($\sim$30%) of known merging clusters hosting double relics have masses $\lesssim 4\times 10^{14}$ M. To investigate the impact of low-velocity shocks on particle acceleration, case studies of relic-hosting low-mass clusters are crucial.
@@
While the population of massive clusters has been extensively characterized in the X-ray, multiwavelength datasets are increasingly detecting low-mass ($\leq 4\times 10^{14}M_{\odot}$) clusters. In this paper, we focus on a detailed investigation of a low-mass cluster PSZ2 G181.06+48.47 in order to uncover its dynamical state and evolution history (Figure 1). The cluster has been little studied in the X-ray wavelength. Originally discovered and confirmed through optical photometric methods at $z\sim 0.24$ (e.g. Koester et al., 2007; Rykoff et al., 2014), the cluster was detected by Planck and presented in the catalog of Sunyaev-Zel’dovich sources (Planck Collaboration et al., 2016) as a low-mass cluster ($M_{\mathrm{SZ}}=(4.2\pm 0.5)\times 10^{14}\,\mathrm{M_{\odot}}$; $r_{\rm 500,SZ}=(1.06\pm 0.05)\,\rm{Mpc}$ [^1] The radius within which the mean density is 500 times the critical density at the cluster redshift.). Double, giant (1.2–1.3 Mpc) radio relics on either side of PSZ2 G181.06+48.47 and a candidate halo at the center were discovered recently (Rajpurohit et al., submitted). The relics were first reported in the LOw Frequency ARray (LOFAR) Two-meter Sky Survey data for 309 Planck SZ-selected galaxy clusters (LoTSS-DR2; Botteon et al., 2022). Both relics are located at very large distances, over $1.2$ Mpc from the cluster center. The cluster, coinciding with the location of diffuse radio sources, strongly suggests a merging system. However, with a Planck detection at a signal-to-noise (S/N) of $\sim 5$, the detailed merger scenario of PSZ2 G181.06+48.47 remained unknown. Nevertheless, the low cluster mass reported by Planck, which implies low particle densities, presents challenges in explaining the positions of relic candidates far apart from each other (Brunetti & Jones, 2014; Kale et al., 2017). PSZ2 G181.06+48.47 thus presents a unique opportunity to study particle acceleration in the low-mass, low-density regime.
@@
We aim to paint a comprehensive picture of PSZ2 G181.06+48.47 through multiwavelength observations spanning the electromagnetic spectrum. In this first paper of the series, we utilize Chandra and XMM-Newton to unambiguously confirm PSZ2 G181.06+48.47 as a merging cluster, measure its detailed properties and characterize its substructure. We search for ICM discontinuities to understand the cluster dynamics. By accurately determining the cluster mass distribution, we also explore viable merger scenarios. Rajpurohit et al. (submitted) presents new multi-band radio observations of the cluster to determine the properties of the diffuse radio sources and derive key insights into the underlying particle acceleration mechanism. Ahn et al. (in prep) leverage a weak-lensing (WL) analysis, complemented by tailored hydrodynamical simulations, to obtain accurate cluster mass estimates, as well as to constrain the merger scenario.
@@ 2.1 XMM-Newton
The data were cleaned for periods of high background induced by solar flares using the tasks mos-filter and pn-filter. We did not find any significant time intervals affected by flares, such that after this cleaning, the net exposure times of MOS1, MOS2, and pn were $\sim$111 ks, $\sim$110 ks, and 92 ks, respectively. Following standard filtering techniques, we kept the single, double, triple, and quadruple events in the MOS data ($\leq$12) and restricted the pn data to single and double events only (pattern $\leq 4$). Additionally, we also excluded all the CCDs in the so-called anomalous state (for more details, see Kuntz & Snowden, 2008). Point-like sources were detected using the task edetect_chain and excluded from the event files. The masked regions can be found in Appendix Figure 12. We used the Filter Wheel Closed (FWC) data, cleaned by applying the same PATTERN selection, flare rejection criteria, and point-source removal as used for the observation events, as instrumental backgrounds event files.
@@ 2.2 Chandra
The smaller point spread function (PSF) of Chandra enables us to cleanly separate point-like and diffuse sources. We identified two extended diffuse clumps located north and south of the cluster, at distances beyond $r_{\rm 500,SZ}$ (see Figure 12). Since these extended sources are located along the merger axis, we examined their potential association with the larger-scale environment of PSZ2 G181.06+48.47. Both the northern (C1) and southern clumps (C2) coincide with background clusters at $z\sim 0.49$ (Wen & Han, 2015) and $z\sim 0.32$ (Zou et al., 2021; Ahn et al., in prep), respectively. Similar clump-like diffuse features observed, for example, in the outskirts of A 133 and A 1795 are predominantly attributed to background clusters that are projected along the line of sight (LOS; Zhu et al., 2023; Kovács et al., 2023). Both clumps were also excluded from any further analysis.
We applied the wavdetect algorithm on the Chandra counts image in the 0.5–2 keV band, adopting a false-positive probability threshold of $10^{-6}$ and wavelet scales of 1, 2, 4, 8, 16, 32 pixels. We excluded the extended clumps (C1 and C2) as well as the point sources detected above. We manually adjusted the wavdetect regions for bright sources to safely exclude the surface brightness wings of their emission. The Chandra flux map, shown in the left panel of Figure 2, was obtained after correcting the NXB-subtracted count image with the combined exposure map and filling the point sources with dmfilth.
@@ 2.3 Ancillary observations
We use multiwavelength data to aid in the interpretation of the X-ray observations, including 144 MHz images from the LOFAR Two-metre Sky Survey (LoTSS; Shimwell et al., 2019; Botteon et al., 2022), optical g, r, and i imaging from the Panoramic Survey Telescope and Rapid Response System survey (Pan-STARRS; Flewelling et al., 2020), and optical spectroscopy from the Sloan Digital Sky Survey Data Release 18 (SDSS DR 18; Almeida et al., 2023). We refer the reader to Rajpurohit et al. (submitted) and Ahn et al. (in prep), respectively, for details on the new multi-band radio and optical observations, respectively on the cluster.
@@ 3.5 Thermodynamic maps
To investigate the temperature, electron density, pressure, and entropy distribution of the merging cores, we follow the method presented in Lovisari et al. (2024), with the regions obtained using the Weighted Voronoi Tessellation (WVT) binning algorithm by Diehl & Statler (2006), which is a generalization of the Cappellari & Copin (2003) Voronoi binning algorithm, by requiring a signal-to-noise S/Nsimilar-$\sim$30. While there are uncertainties associated with any choice, in deriving the density map, we assume that each emitting region has homogeneous properties and a volume of $V=2A\sqrt{r_{500,P}^{2}-X^{2}-Y^{2}}$, where $A$ is the area of the region, and $X$ and $Y$ are the projected distances in the east-west and north-south directions, respectively.
@@ 4.1 Morphological features
As shown in Figure 2, PSZ2 G181.06+48.47 has an X-ray morphology elongated along the northeast-southwest direction, indicating a cluster merger process is under way along this axis. The cluster displays two X-ray peaks (highlighted with magenta contours in Figure 1, as well as in the GGM-filtered images in Figure 4) separated by $\sim 370$ kpc and connected through a bridge of emission, likely tracing two distinct subclusters observed post-core passage. The concentration ($c$) and centroid-shift ($w$) are among the most robust and widely used parameters to probe the dynamical state of clusters. The concentration parameter is defined as the ratio of the X-ray integrated fluxes in the cluster core and the surrounding outer region of a galaxy cluster. The centroid-shift is defined as the standard deviation of the projected distance between the X-ray peak and the cluster’s center. The concentration $c=0.087\pm 0.001$ and centroid-shift $w=0.053\pm 0.001$ parameters place PSZ2 G181.06+48.47 among the most disturbed clusters: out of 120 clusters studied by Lovisari et al. (2017) only five had both a lower concentration and a higher centroid-shift.
Both subclusters are compact, with the southern subcluster being brighter (see Figure 2 and 4). The morphology is reminiscent of typical binary, post-core-passage merging clusters, such as CIZA J2242.8+5301, PSZ2 G096.88+24.18, or RXC J1314.4-2515 (Golovich et al., 2019). The southern subcluster is about 1.3 times more luminous in the 0.5–2 keV band than the northern one, supporting a major merger scenario. The spectroscopically-confirmed SDSS DR18 galaxies have an elongated distribution, which matches the merger axis inferred from the X-ray morphology. By cross-matching with the Pan-STARRS imaging and SDSS DR18 spectroscopy (see Figure 1), we identify no magnitude gap ($\Delta m_{\rm Rband}=0$) between the two brightest cluster members. The positions of both BCGs are consistent with the locations of their corresponding X-ray peaks (middle panel in Figure 1).
@@
We also detect a clump of emission, $\sim 65$ kpc in extent, toward the east of the northern subcluster. The clump is clearly detected as a relatively compact peak both in the X-ray maps (Figure 2) and in the unsharp-masked and GGM-filtered images (Figure 4), and is not associated with any point source or background cluster. While the emission could be associated with a subcluster, given that the clump is offset from any significant dark matter clump (see WL map in Ahn et al., in prep), its nature is uncertain.

### v3

@@ Abstract
Relics are diffuse, highly-polarized radio sources that trace merger-driven shocks at the periphery of merging galaxy clusters. The LOFAR survey recently discovered a rare example of double relics in the low-mass cluster PSZ2 G181.06+48.47. Through a detailed exploration of new Chandra and XMM-Newton observations, we reveal that PSZ2 G181.06+48.47 has a lower mass ($M_{\rm 500,X}=2.32^{+0.29}_{-0.25}\times 10^{14}\;{\rm M_{\odot}}$) than previously thought. Despite its cool global temperature of $kT_{500}=3.62^{+0.15}_{-0.07}$ keV, PSZ2 G181.06+48.47 is one of the most disturbed clusters in the Planck sample, with a complex morphological and thermodynamic structure. We discover a set of three discontinuities within $<500$ kpc of the cluster center, and, from a surface brightness analysis, place $5\sigma$ upper limits of $\mathcal{M}_{NE}<1.43$ and $\mathcal{M}_{SW}<1.57$ for any shock associated with the relic locations. We also revise established scaling relations for double radio-relics by adding 12 new systems not included in previous work. The PSZ2 G181.06+48.47 relics have the widest separation (scaled for $r_{500}$) of all known double-relic systems. The exceptional distance from the cluster center ($>r_{200}$), indicates the relics may be associated with shocks in the “run-away” phase. We propose that this late-stage, post-apocenter merger is captured as the two subclusters with a mass ratio of 1.2–1.4 fall back into each other. The outer relic shocks were likely produced at the first core passage, while the inner discontinuities are associated with the second infall.
@@
journal: ApJfacilities: XMM, CXO, LOFAR, SDSS, Pan-STARRSsoftware: pyproffit (Eckert et al., 2020), Astropy (Astropy Collaboration et al., 2013, 2018), APLpy (Robitaille & Bressert, 2012), Matplotlib (Hunter, 2007), TOPCAT (Taylor, 2005), DS9 (Joye & Mandel, 2003), CARTA (Comrie et al., 2021), CIAO (Fruscione et al., 2006), SAS
@@ 1 Introduction
Thermal and non-thermal components of the ICM such as shocks, cold fronts, turbulence, magnetic fields, and cosmic ray particles make clusters unique laboratories to study plasma physics (e.g. Sarazin, 1986, 2004; van Weeren et al., 2019). Thanks to the combination of sub-arcsecond angular resolution (0.5″, with Chandra) and sensitivity to diffuse extended thermal emission achievable by X-ray telescopes, the impact that merger events have on the plasma atmospheres and substructures can be directly mapped, for example, through sharp X-ray surface brightness edges, such as shocks and cold fronts (e.g. Markevitch et al., 2002, 2005; Markevitch & Vikhlinin, 2007; Owers et al., 2009; Ogrean & Brüggen, 2013; Eckert et al., 2016; Akamatsu et al., 2017; Zhang et al., 2021b; Diwanji et al., 2024). Unlike shocks, where the gas density and temperature are higher in the downstream compared to the upstream, cold fronts are formed at the boundary of gas clouds moving through a hotter and rarefied surrounding medium (see the review by Markevitch & Vikhlinin, 2007). Cold fronts have been detected in both merging and relaxed clusters (e.g., Botteon et al., 2018; Ghizzardi et al., 2010; Markevitch & Vikhlinin, 2007; Paterno-Mahler et al., 2013).
Figure 1: Multiwavelength view of PSZ2 G181.06+48.47, showing the central $15\arcmin\times 14\arcmin$ area. Left: 0.5–2 keV XMM-Newton image of the cluster, with LOFAR 140 MHz radio contours drawn at $[1,2,4,8...]\times 3.0\sigma_{\rm rms}$, unveiling the cluster as a clear merger with an elongated, double-peaked X-ray morphology. Center: RGB Pan-STARRS image, with radio (green) and XMM-Newton (magenta, drawn at $[2,4,6,8...]\times 10^{-8}\,{\rm cts/s/pixel}$) contours overplotted. Saffron yellow squares mark spectroscopically-confirmed cluster members, including two BCGs roughly following the XMM-Newton X-ray peaks. The white plus marks the center of the system. Right: The 140 MHz LOFAR image at 6″ resolution reveals two spectacular diffuse radio sources on either side of the XMM-Newton X-ray detection of the cluster ICM (contours in magenta). The white, dashed circle represents $r_{\rm 500,SZ}=1.06$ Mpc.
@@
In rare cases, merging systems can exhibit two relics on opposite sides of the cluster. To our knowledge, only 30 examples of merging clusters hosting double radio relics have been discovered so far (de Gasperin et al., 2014; van Weeren et al., 2019; Botteon et al., 2020a; Jones et al., 2023). For example, only about 1/3 of simulated radio relic clusters may host candidate double relics of relatively similar luminosity (luminosity ratio $>0.3$, Lee et al., in prep). More importantly, double relics are more likely to be observed in simple binary cluster mergers with a favorable orientation close to the plane-of-the-sky (Golovich et al., 2019), providing a unique laboratory to study the shocks with less interference by projection effects.
@@
While many X-ray shocks associated with radio relics have been successfully studied, shocks created in low-mass clusters with lower ICM sound speeds and, hence, slower shock velocities remain under-explored (Eckert et al., 2016; Akamatsu et al., 2017; Urdampilleta et al., 2018). Such shocks probe a key region of parameter space and are critical to constraining shock acceleration models. For example, standard diffusive shock acceleration (DSA, e.g. Bell, 1978; Drury, 1983; Blandford & Eichler, 1987), where particles from the thermal pool are directly accelerated by the merger-induced shock front, has been proposed as an effective mechanism for the origin of radio relics (for details see Brunetti & Jones, 2014). For low Mach number shocks such as those in clusters ($\mathcal{M}<2$; Ryu et al., 2019), DSA is not sufficient to explain the observed bright emission associated with large radio relics due to the low acceleration efficiency (e.g. Botteon et al., 2020a). Only a small fraction ($\sim$30%) of known merging clusters hosting double relics have masses $\lesssim 4\times 10^{14}$ M. To investigate the impact of low-velocity shocks on particle acceleration, case studies of relic-hosting low-mass clusters are crucial.
@@
While the population of massive clusters has been extensively characterized in the X-rays, multiwavelength datasets are increasingly detecting low-mass ($\leq 4\times 10^{14}M_{\odot}$) clusters. In this paper, we focus on a detailed investigation of a low-mass cluster PSZ2 G181.06+48.47 in order to uncover its dynamical state and evolution history (Figure 1). The cluster has been little studied in the X-ray band. Originally discovered and confirmed through optical photometric methods at $z\sim 0.24$ (e.g. Koester et al., 2007; Rykoff et al., 2014), the cluster was detected by Planck and presented in the catalog of Sunyaev-Zel’dovich sources (Planck Collaboration et al., 2016) as a low-mass cluster ($M_{\mathrm{SZ}}=(4.2\pm 0.5)\times 10^{14}\,\mathrm{M_{\odot}}$; $r_{\rm 500,SZ}=(1.06\pm 0.05)\,\rm{Mpc}$ [^1] The radius within which the mean density is 500 times the critical density at the cluster redshift.). Double, giant (1.2–1.3 Mpc) radio relics on either side of PSZ2 G181.06+48.47 and a candidate halo at the center were discovered recently (Rajpurohit et al., 2025). The relics were first reported in the LOw Frequency ARray (LOFAR) Two-meter Sky Survey data for 309 Planck SZ-selected galaxy clusters (LoTSS-DR2; Botteon et al., 2022). Both relics are located at very large distances, over $1.2$ Mpc from the cluster center. The cluster, coinciding with the location of diffuse radio sources, appears to be a merging system. However, with a Planck detection at a signal-to-noise (S/N) of $\sim 5$, the detailed merger scenario of PSZ2 G181.06+48.47 remained unknown. Nevertheless, the low cluster mass reported by Planck, which implies low particle densities, presents challenges in explaining the positions of relic candidates far apart from each other (Brunetti & Jones, 2014; Kale et al., 2017). PSZ2 G181.06+48.47 thus presents a unique opportunity to study particle acceleration in the low-mass, low-density regime.
@@
We aim to paint a comprehensive picture of PSZ2 G181.06+48.47 through multiwavelength observations spanning the electromagnetic spectrum. In this first paper of the series, we utilize Chandra and XMM-Newton to unambiguously confirm PSZ2 G181.06+48.47 as a merging cluster, measure its detailed properties and characterize its substructure. We search for ICM discontinuities to understand the cluster dynamics. By accurately determining the cluster mass distribution, we also explore viable merger scenarios. Rajpurohit et al. (2025) presents new multi-band radio observations of the cluster to determine the properties of the diffuse radio sources and derive key insights into the underlying particle acceleration mechanism. Ahn et al. (2025) leverage a weak-lensing (WL) analysis, complemented by tailored hydrodynamical simulations, to obtain accurate cluster mass estimates, as well as to constrain the merger scenario.
@@ 2.1 XMM-Newton
The data were cleaned for periods of high background using the tasks mos-filter and pn-filter. We did not find any significant time intervals affected by flares, such that after this cleaning, the net exposure times of MOS1, MOS2, and pn were $\sim$111 ks, $\sim$110 ks, and 92 ks, respectively. Following standard filtering techniques, we kept the single, double, triple, and quadruple events in the MOS data ($\leq$12) and restricted the pn data to single and double events only (pattern $\leq 4$). Additionally, we also excluded all the CCDs in the so-called anomalous state (for more details, see Kuntz & Snowden, 2008). Point-like sources were detected using the task edetect_chain and excluded from the event files (see Figure 4). We used the Filter Wheel Closed (FWC) data, properly rescaled to match our observation (see Section 3.3), as instrumental backgrounds event files.
@@ 2.2 Chandra
The smaller point spread function (PSF) of Chandra enables us to cleanly separate point-like and diffuse sources. We identified two extended diffuse clumps located north and south of the cluster, at distances beyond $r_{\rm 500,SZ}$. Since these extended sources are located along the merger axis, we examined their potential association with the larger-scale environment of PSZ2 G181.06+48.47. Both the northern and southern clumps coincide with background clusters at $z\sim 0.49$ (Wen & Han, 2015) and $z\sim 0.32$ (Zou et al., 2021; Ahn et al., 2025), respectively. Similar clump-like diffuse features observed, for example, in the outskirts of A 133 and A 1795 are predominantly attributed to background clusters that are projected along the line of sight (LOS; Zhu et al., 2023; Kovács et al., 2023). Both clumps were also excluded from any further analysis.
We applied the wavdetect algorithm on the Chandra counts image in the 0.5–2 keV band, adopting a false-positive probability threshold of $10^{-6}$ and wavelet scales of 1, 2, 4, 8, 16, 32 pixels. We excluded the extended clumps as well as the point sources detected above. We manually adjusted the wavdetect regions for bright sources to safely exclude the surface brightness wings of their emission. The Chandra flux map, shown in the left panel of Figure 2, was obtained after correcting the NXB-subtracted count image with the combined exposure map and filling the point sources with dmfilth.
@@ 2.3 Ancillary observations
We use multiwavelength data to aid in the interpretation of the X-ray observations, including 144 MHz images from the LOFAR Two-metre Sky Survey (LoTSS; Shimwell et al., 2019; Botteon et al., 2022), optical g, r, and i imaging from the Panoramic Survey Telescope and Rapid Response System survey (Pan-STARRS; Flewelling et al., 2020), and optical spectroscopy from the Sloan Digital Sky Survey Data Release 18 (SDSS DR 18; Almeida et al., 2023). We refer the reader to Rajpurohit et al. (2025) and Ahn et al. (2025), respectively, for details on the new multi-band radio and optical observations, respectively on the cluster.
@@ 3.5 Thermodynamic maps
To investigate the temperature, electron density, pressure, and entropy distribution of the merging cores, we follow the method presented in Lovisari et al. (2024), with the regions obtained using the Weighted Voronoi Tessellation (WVT) binning algorithm by Diehl & Statler (2006), which is a generalization of the Cappellari & Copin (2003) Voronoi binning algorithm, by requiring a signal-to-noise S/Nsimilar-$\sim$30. While there are uncertainties associated with any choice, in deriving the density map, we assume that each emitting region has homogeneous properties and a volume of $V=2A\sqrt{r_{500,P}^{2}-X^{2}-Y^{2}}$, where $A$ is the area of the region, and $X$ and $Y$ are the projected distances in the east-west and north-south directions from the cluster center, respectively.
@@ 4.1 Morphological features
As shown in Figure 2, PSZ2 G181.06+48.47 has an X-ray morphology elongated along the northeast-southwest direction, indicating a cluster merger process is under way along this axis. The cluster displays two X-ray peaks (highlighted with magenta contours in Figure 1, as well as in the GGM-filtered images in Figure 3) separated by $\sim 370$ kpc and connected through a bridge of emission, likely tracing two distinct subclusters observed post-core passage. The concentration ($c$) and centroid-shift ($w$) are among the most robust and widely used parameters to probe the dynamical state of clusters. The concentration parameter is defined as the ratio of the X-ray integrated fluxes in the cluster core and the surrounding outer region of a galaxy cluster. The centroid shift is measured as the standard deviation of the projected separation between the X-ray peak and the centroid (within an aperture radius 500 kpc). It is computed in a series of circular apertures centered on the cluster X-ray peak. The concentration $c=0.087\pm 0.001$ and centroid-shift $w=0.053\pm 0.001$ parameters place PSZ2 G181.06+48.47 among the most disturbed clusters: out of 120 clusters studied by Lovisari et al. (2017) only five had both a lower concentration and a higher centroid-shift.
Both subclusters are compact, with the southern subcluster being brighter (see Figure 2 and 3). The morphology is reminiscent of typical binary, post-core-passage merging clusters, such as CIZA J2242.8+5301, PSZ2 G096.88+24.18, or RXC J1314.4-2515 (Golovich et al., 2019). The southern subcluster is about 1.3 times more luminous in the 0.5–2 keV band than the northern one, supporting a major merger scenario. The spectroscopically-confirmed SDSS DR18 galaxies have an elongated distribution, which matches the merger axis inferred from the X-ray morphology. By cross-matching with the Pan-STARRS imaging and SDSS DR18 spectroscopy (see Figure 1), we identify no magnitude gap ($\Delta m_{\rm Rband}=0$) between the two brightest cluster members. The positions of both BCGs are consistent with the locations of their corresponding X-ray peaks (middle panel in Figure 1).
@@
We also detect a clump of emission, $\sim 65$ kpc in extent, toward the east of the northern subcluster. The clump is clearly detected as a relatively compact peak both in the X-ray maps (Figure 2) and in the unsharp-masked and GGM-filtered images (Figure 3), and is not associated with any point source or background cluster. While the emission could be associated with a subcluster, given that the clump is offset from any significant dark matter clump (see WL map in Ahn et al., 2025), its nature is uncertain.